OpenStudy (cenaida):

Write the equation of the line with the given slope, 2 and y-intercept, 1/2. A. y = (1/2)x - 2 B. y = (1/2)x + 2 C. y = 2x - (1/2) D. y = 2x + (1/2)

OpenStudy (owlcoffee):

Remember: The equation of the line has this form: \[y=mx+b\] Where "m" is the slope and "b" is the y-intersection.

OpenStudy (anonymous):

Just write this Modom : 

OpenStudy (cenaida):

I don't understand

OpenStudy (anonymous):

In this get 2___>slope 1/2___>y-intercept

OpenStudy (cenaida):

so which one is the answer?

OpenStudy (owlcoffee):

We won't give it to you, we'll only give you a push :). Wich one do you think it is?

OpenStudy (anonymous):

@Owlcoffee:We help every one if want ....

OpenStudy (owlcoffee):

I know, but the idea is that she learns because we won't be all the time answering :(

OpenStudy (cenaida):

is it B ?

OpenStudy (owlcoffee):

ooh... no, remember the placing of the slope and the y-intersection :)

OpenStudy (anonymous):

Not realy ! You have a small wrong ! 2__>slope 1/2___>intercep Change the place of them :)

OpenStudy (cenaida):

oooooh so its D!

OpenStudy (anonymous):

Yea Modom ! Did u get it ?:)

OpenStudy (cenaida):

yess so is tis one D too? Write the equation of the line with the given slope, 1/2 and y-intercept, 7. A. y = (1/2)x - 7 B. y = (1/2)x + 7 C. y = 7x - ½ D. y = 7x + ½

OpenStudy (anonymous):

Answer this Modom ! You can ..:)

OpenStudy (anonymous):

y=2x+1/2